### v3

# ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images

Abstract
In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting conditions and obstructions (e.g., handwear). The benchmark includes a multi-view and multi-spectral dataset collected from 28 subjects performing hand-object and hand-virtual interactions under diverse scenarios, accurately annotated with 3D hand poses through an automated process. We introduce a new baseline method, TherFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery. Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
1 Introduction
Egocentric 3D hand pose estimation is critically important for interpreting hand gestures across various applications, ranging from extended reality (XR) [1, 2, 3, 4], to human-robot interaction [5, 6, 7], and to imitation learning [8, 9, 10]. Its importance has been magnified with the advent of advanced XR headsets such as the Meta Quest series [11] and Apple Vision Pro [12], where it serves as a cornerstone for spatial interaction and immersive digital experiences.
While current research of hand pose estimation primarily focuses on RGB image-based methods [13, 14, 15, 16, 17], these approaches are particularly vulnerable to issues related to lighting variation and occlusions caused by handwear, e.g., gloves or large jewellery [18, 19]. These challenges underscore the imperative for robust egocentric 3D hand pose estimation capable of performing reliably in a variety of common yet complex daily scenarios. The prevailing approach to facilitate robust hand pose estimation in low-light conditions utilizes near infrared (NIR) cameras paired with active NIR emitters. This technology, invisible to the human eye, leverages active NIR emitter-receiver configurations for depth estimation through time-of-flight (ToF) or structured lighting. Nevertheless, active NIR systems are more power-intensive compared to passive sensing technologies [20, 21] and are prone to interference from external NIR sources, such as sunlight [22] and other NIR-equipped devices [23]. Consequently, these vulnerabilities restrict the effectiveness of hand pose estimation under bright daylight conditions and in situations where multiple augmented reality (AR) or virtual reality (VR) systems are used for collaborative works.
In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [24]. This unique attribute of thermal imaging introduces several benefits for hand pose estimation. Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability. Moreover, thermal cameras can detect hands even under handwear such as gloves by identifying heat transmission patterns. This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions. To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab. 1). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [25] (cfFig. 3).
Together with the multi-spectral dataset, we introduce a new baseline method named TherFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cfFig. 4). This approach is notable for its two consecutive transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency. Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cfTab. 2). We then benchmark TherFormer against leading methods (cfTab. 3) and compare the performance of various spectral images (cfTab. 4 and Fig. 6). The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques. Our main contributions are summarized as follows:
  • $\bullet$
    We introduce the first-of-its-kind benchmark, dubbed ThermoHands, to investigate the potential of thermal imaging for egocentric 3D hand pose estimation.
  • $\bullet$
    We collected a diverse dataset comprising approximately 96,000 synchronized multi-spectral, multi-view images capturing hand-object and hand-virtual interactions from 28 participants across various environments. This dataset is enriched with 3D hand pose ground truths through an innovative automatic annotation process.
  • $\bullet$
    We introduce a new baseline method, termed TherFormer, and implement state-of-the-art image-based methods on our dataset for benchmarking.
  • $\bullet$
    Based on the ThermoHands benchmark, we conduct comprehensive experiments and analysis on TherFormer and state-of-the-art methods.
  • $\bullet$
    We will release our dataset, code and models and maintain the benchmark to serve as a new challenge in 3D hand pose estimation.
2 Related Works
2.1 3D Hand Pose Datasets
Datasets with 3D hand pose annotations are imperative for training and evaluating ad-hoc models. Existing datasets, according to their approaches of annotation acquisition, can be summarized as four types in general, i.e., marker-based [26, 27, 28, 29], synthetic [30, 31, 32, 33, 34], manual [30, 35] or hybrid [36, 37, 38, 39, 40], and automatic [41, 42, 43, 44] annotated datasets. Marker-based approaches, using magnetic sensor [26, 27] or Mocap markers [28, 29], can alter and induce bias to the hand appearance. Synthetic data [30, 31, 33, 34, 32] suffers from the sim2real gap in terms of hand motion and texture features. Introducing human annotators to fully [30, 35] or partly [36, 37, 38, 39, 40] annotate 2D/3D keypoints circumvents the issues above, but it either limits the scale of datasets or manifests costly and laborious in practice. Most similar to ours, some datasets adopt fully automatic pipelines to obtain 3D hand pose annotations [41, 42, 43], which leverage pre-trained models (e.g. OpenPose [45]) to infer the prior hand information and rely on optimization to fit the MANO hand model [25].
Despite the existing progress, previous datasets only provide depth [26, 32], RGB images [33, 38, 37, 40] or both of them [27, 28, 31, 34, 30, 35, 36, 39, 41, 42, 43] as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation. ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images. Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications.
Figure 1: Data capture setup with the HMSP and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
2.2 Image-based 3D Hand Pose Estimation
As a key computer vision task, 3D hand pose estimation from images is highly demanded by applications like XR [1, 2, 3, 4], human-robot interaction [5, 6, 7] and imitation learning [8, 9, 10]. Therefore, this field has been extensively explored in previous arts that uses single RGB [46, 47, 48, 49, 50, 13, 38, 14, 51, 52, 53, 31, 54, 55, 56, 57] or depth [58, 59, 60, 61, 62] image as input. These methods can be roughly categorized into two fashions, i.e., model-based and model-free methods. Model-based methods [14, 51, 52, 53, 31, 54, 55, 56, 57] utilize the prior knowledge of the MANO hand model [25] by estimating its shape and pose parameters, while model-free methods [46, 47, 48, 49, 50, 13, 38, 58, 59, 60, 61, 62] learn the direct regression of 3D hand joints or vertices coordinates. Recently there has been growing interest in leveraging the temporal supervision [63, 54, 64, 65, 66] or leveraging sequential images as input [15, 16, 17, 67, 68, 69] for 3D hand pose estimation. In this study, we evaluate existing methods and our baseline method in both single image-based and video-based problem settings, respectively. Apart from the previous approaches, we investigate the potential of thermal imagery for tackling various challenges in 3D hand pose estimation.
2.3 Thermal Computer Vision
Thermal cameras achieve imaging by capturing the radiation emitted in the LWIR spectrum and deducing the temperature distribution on the surfaces [24]. Leveraging its robustness to variable illumination and unique temperature information, numerous efforts have been made to address various computer vision tasks, including super-resolution [70, 71, 72], human detection [73, 74, 75], action recognition [76, 77] and pose estimation [78, 79, 80], semantic segmentation [81, 82, 83, 84], depth estimation [85, 86, 87, 88], visual(-inertial) odometry/SLAM [89, 90, 91, 92], 3D reconstruction [93, 94, 95], etc. In this work, we focus on 3D hand pose estimation, which is an under-exploited task based on thermal images.
3 The ThermoHands Benchmark
3.1 Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab. 1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from School of Informatics, University of Edinburgh, and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the normal office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
Normal office (Main) Other settings Total
Setting train val test sum darkness sun glare gloves kitchen
#frames 47,436 12,914 24,002 84,352 3,188 2,508 3,068 2,808 95,924
#seqs 172 43 86 301 12 12 12 14 352
#subjects 16 4 8 28 1 1 1 2 -
Table 1: Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
Figure 2: Thermal calibration chessboard containing a black base board and multiple removable white cubes (a). By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
Sensor Platforms. The HMSP is mounted with an Intel RealSense L515 LiDAR depth camera [20] streaming egocentric RGB, depth and NIR images, and a FLIR Boson 640 thermal camera [96] receiving the LWIR to obtain the thermal images. To better simulate XR devices, the HMSP is modularly designed as two components: a base and a sensor board with a fixed 30-degree downward tilt. An extra exocentric platform equipped with an Intel RealSense D455 [21] is leveraged to support multi-view annotation (cf. Sec. 3.2) as well as provide the RGB-D image data from the third-person viewpoint. As exhibited in Fig. 1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters [20, 21]. In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
Synchronization. We use a single PC to simultaneously gather data streams from two sensor platforms, ensuring the synchronization of their timestamps. After collection, we synchronize six types of images, each with distinct frame rates, w.r.t. the timestamps of thermal images (8.5fps), thereby generating well-aligned multi-spectral, multi-view data samples as our released data.
Calibration. For accuracy, we use factory-calibrated intrinsic parameters from the D455 and L515 cameras [20, 21]. To better calibrate the parameters of the thermal camera, we self-design a modular calibration chessboard as shown in Fig. 2. Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig. 2. In this way, we can simultaneously calibrate the intrinsic parameter of the L515 thermal camera and its extrinsic parameter w.r.t. the D455 RGB-D camera. To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table. At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP problem [97]. For the subsequent frames, as the exocentric platform keeps stationary, we only track the the motion of the ego-head using KISS-ICP [98] odometry method, which takes the point clouds converted from the egocentric depth image as input.
Dataset Statistics. As seen in Tab. 1, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cfFig. 1) and 352 independent sequences in total. The main part is collected under the normal office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions. performs 7 hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands. This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2. We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen. Please refer to the supplementary for more dataset details.
3.2 Hand Pose Annotation
To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [41, 42, 43], to obtain the 3D hand pose ground truth for our dataset. In particular, we use the MANO statistical hand model [25] to represent 3D hand pose. The MANO model parameterizes the hand mesh vertices into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$. The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$ (1)
where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 3. For more explanation of error terms and implementation details of our dataset annotation process, please refer to our supplementary document.
Figure 3: Automatic annotation pipeline of 3D hand pose. We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands [99] and SAM [100]. Various error terms are formulated to optimize the MANO parameters.
3.3 TherFormer: A Baseline Method
For our benchmark, we setup a baseline method, dubbed TherFormer, for thermal image-based 3D hand pose estimation. As exhibited in Fig. 4, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient. Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images. However, our primary objective is to establish a baseline method for future research, rather than focusing exclusively on methods for thermal images.
Figure 4: Overall Framework of TherFormer. Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction. Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
Problem Definition. We consider two problem settings: single image-based and video-based egocentric 3D hand pose estimation for our benchmark. In the former setting, we aim to estimate the 3D joint positions $\mathcal{J}_{t}$ for two hands given the single thermal image $\mathcal{I}_{t}$ captured for the $t$-th frame. For the video-based one, our input is a sequence of thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ and we estimate the per-frame 3D hand joint positions $\mathcal{J}_{i}$ together. Without losing the generality, here we illustrate our network architecture for the video-based setting, i.e., TherFormer-V. In practice, our network can be flexibly adapted to the single image-based version (i.e., TherFormer-S) by setting $T=1$.
Mask-guided Spatial Transformer. Human hands are highly articulated objects that can adopt a wide variety of poses, often against complex backgrounds. We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction. Given backbone features, we first utilize a mask head to estimate the binary hand mask in the thermal image. Then, we leverage the deformable self-attention [101] to refine the hand spatial features under the guidance of the estimated hand mask. Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations. In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter. Lastly, we reduce the spatial dimensions of the spatial features with a series of convolutions layers for efficiency.
Temporal Transformer. Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities. To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention [102] to explicitly attend to the feature vector of every frame. The output is frame-wise spatio-temporal feature embeddings. Note that the temporal self-attention degrades to an MLP for single-image based setting.
In the pose head, we use the MLP to project the embeddings to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$. We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cfSec. 3.2). For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately. Please see more details of TherFormer in our supplementary document.
4 Experiment
4.1 Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec. 3.2) and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames. To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D joint positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
As shown in Tab. 2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [43, 41, 37]. The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture. We also observe that only combining $\mathcal{L}_{mask}$ and $\mathcal{L}_{j2d}$ can already provide a plausible accuracy since they fit the projection of the 3D hand pose to two heterogeneous views. $\mathcal{L}_{mesh}$, though it hardly improves the joint accuracy, can result in more natural hand mesh. Adding $\mathcal{L}_{j3d}$ further refines the joints as it induces the explicit constraint to their positions. We showcase some annotation examples in Fig. 5. As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subjects’ hand color and shape.
Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
Ego-view optimization Multi-view optimization
Errors $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}$ $\mathcal{L}_{mask}+\mathcal{L}_{j2d}+\mathcal{L}_{mesh}+\mathcal{L}_{j3d}$
mean (std) 37.29 ($\pm$ 18.02) 7.03 ($\pm$ 2.57) 8.13 ($\pm$ 0.57) 1.29 ($\pm$ 0.43) 1.28 ($\pm$ 0.43) 1.01 ($\pm$ 0.34)
Table 2: Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
4.2 Experiment Setup
Dataset Preparation. We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments. We annotate the main part of our dataset (cfTab. 1) automatically following Sec. 3.2, of which the training and validation sets serve as the foundation for the training of all network models.
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT [15] and A2J-Transformer [46], and reproduced them on our dataset for experiments. HTT [15] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [46] only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines. We exclude the additional action block used by HTT [15], focusing solely on hand pose estimation. We adjusted the anchor initialization phase of A2J-Transformer [46] to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
Evaluation Metrics. We evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) [34], in both camera space and root-aligned (RA) space [38]. For RA space, we align the estimated wrist with its groundtruth position before measurement. For PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
4.3 Thermal Image-based 3D Hand Pose Estimation
Main Results. To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods [15, 46] on the main testing set, as shown in Tab. 3. TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT [15] given the same sequential images as input. Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec. 3.3). A performance gap can be observed between single image-based and video-based settings for both HTT [15] and TherFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices. Moreover, TherFormer-V also improves the performance over HTT [15] for other spectra (cfTab. 4), proving its adaptability to different inputs.
Method Input MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC-RA $\uparrow$ fps $\uparrow$
(a) A2J-Transformer [46] Single 51.68 0.474 20.76 0.603 34
(b) HTT [15] Single 49.09 0.489 20.69 0.599 211
(c) TherFormer-S Single 44.64 0.539 18.34 0.643 136
(d) (c) w/o spatial transformer Single 48.79 0.491 20.15 0.609 174
(e) (c) w/o mask guidance Single 48.83 0.494 18.89 0.625 141
(f) HTT [15] Sequence 47.07 0.512 17.49 0.659 129
(g) TherFormer-V Sequence 43.36 0.549 17.36 0.661 52
Table 3: Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
Ablation Studies. We ablate our proposed mask-guided attention mechanism and the entire spatial transformer (cfSec. 3.3) to observe their impact. As can be seen in Tab. 3, the mask-guided attention mechanism notably contributes to TherFormer’s performance growth (row (c) vs. (e)), demonstrating its effectiveness in mitigating the effect of background clutter. Other components in the spatial transformer bring a large performance gain in the RA space (row (e) vs. (d)), confirming the importance of spatial feature enhancement for fine-grained hand joints localization.
4.4 Comparison between Spectrum
4.4.1 Qualitative Results under Challenging Conditions
To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf. Tab. 1). Our automatic annotation pipeline (cf. Sec. 3.2) becomes infeasible since hands appear corrupted in either RGB or depth images. Therefore, we conduct a qualitative analysis of their performance and present some representative examples in Fig. 6. We use the same model, i.e., HTT [15], trained on different spectra for a fair comparison. Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
RGB vs. Thermal. As can be seen, RGB-based methods fail with gloves wearing (cfFig. 6 (a-b)) and in darkness (cfFig. 6 (c-d)). Gloves change how human hands look and hide their natural colors and textures. Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process. Contrary to RGB techniques, as shown in Fig. 6 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
NIR, Depth vs. Thermal. NIR sensors are significantly disrupted by strong sunlight, affecting both NIR imaging and depth map creation, as shown in Fig. 6 (e-f). Conversely, thermal imaging is immune to sunlight and outdoor conditions. The temperature difference between hands and their surroundings in the thermal spectrum facilitates effortless identification of the hands, leaving the thermal-based estimator unaffected. The thermal camera’s ability to consistently capture hand features in diverse lighting conditions positions it as a suitable option for future XR applications.
Figure 6: Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
4.4.2 Quantitative Results under Normal Conditions
To validate the versatility of thermal-based methods, we also evaluate their performance compared to other spectra (i.e., RGB, depth, NIR and thermal) under normal conditions for general-purpose use cases. We apply different spectral images as input to multiple baselines methods and report the results in Tab. 4. Not surprisingly, depth image-based methods yield the best performance on average, as they directly utilize the depth information to provide a detailed 3D structure of the hand. The NIR spectrum shows a marginal decrease in performance compared to depth but outperforms both RGB and thermal spectra. This can be attributed to NIR camera’s active sensing ability, which leads to more consistent and reliable imaging regardless of the variability of external illuminations (vs. RGB) and temperature (vs. thermal). However, active NIR sensors are prone to interference from external NIR sources like sun glare (cfFig. 6). Thermal-based methods, despite using single-channel heat capture, achieves comparable performance to those using RGB images that contain rich color and textures. Such results suggest that thermal images can not only serve as supplements in challenging cases but also can be a viable alternative to other spectra in normal conditions for hand pose estimation.
HTT (Sequence) TherFormer-V TherFormer-S Average
Spectrum MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$ MEPE (mm) $\downarrow$ AUC $\uparrow$
RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.80 0.538
Depth 41.62 0.559 39.70 0.581 39.84 0.579 40.39 0.573
NIR 41.57 0.562 40.79 0.575 40.98 0.575 41.11 0.571
Thermal 47.07 0.512 43.36 0.549 44.64 0.539 45.06 0.533
Table 4: Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
5 Limitation and Future Work
Despite the advancements, we acknowledge certain limitations here guiding our future work. First, the scale of our current dataset is relatively small when compared to some datasets [29, 40, 39, 38]. We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions. Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands. Further efforts will be annotating the fine-grained hand action splits and contact heatmap for hand grasp [42]. Lastly, we currently rely on qualitative analysis for challenging conditions when our automatic annotation method becomes unreliable. In the future, we will leverage unaffected spectral images in individual cases to manually annotate the 3D hand pose.
6 Conclusion
This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images. It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships. We demonstrate near 1cm annotation accuracy, show that TherFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios. We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.
Acknowledgments and Disclosure of Funding
This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1), and partially supported by Cisco Research.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] We elaborate our contributions as a list in the introduction section and briefly explain them in the abstract.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Sec. 5
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] There is no societal impact of this work.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We have reviewed the ethics review guidelines and make sure we obey all rules for this research.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] This work does not contribute to the theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] This work does not contribute to the theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [No] We will release our code upon the acceptance of this paper.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We describe them in our main paper and supplementary document.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] In line with the state-of-the-art 3D hand pose estimation works, we do not report error bars for our experiment results.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We include the type of GPUs in the main paper.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We use the offical code of HTT [15] and A2J-transformer [46] to reproduce their results on our dataset, which have been cited.
    2. (b)
      Did you mention the license of the assets? [Yes] We mention the license of the offical codes of HTT and A2J-Transformer in the supplementary document.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We include the datasheet and the download URL of our dataset in the supplemental document.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] We claim in the footnote that participant consent forms were signed before the collection.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We discuss in the supplementary document.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We do not give written instructions to our participants.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] There is no potential participant risks.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] Our participants are volunteer.